### v1

# Redefining Fitness: Evolution as a Dynamic Learning Process

Abstract
Evolution is the process of optimal adaptation of biological populations to their living environments. This is expressed via the concept of fitness, defined as relative reproductive success. However, it has been pointed out that this definition is incomplete and logically circular. To address this issue, several authors have called for new ways to specify fitness explicitly in terms of the relationship between phenotypes and their environment. Here, we show that fitness, defined as the likelihood function that follows from mapping population dynamics to Bayesian learning, provides a general solution to this problem. We show how probabilistic models of fitness can easily be constructed in this way, and how their averages acquire meaning as information. We also show how this approach leads to powerful tools to analyze challenging problems of evolution in variable environments, game theory, and selection in group-structured populations. The approach is general and creates an explicit bridge between population dynamics, statistical learning theory, and emerging models of artificial intelligence.
keywords:
Adaptation, Statistical Mechanics, Optimization, Game Theory, Phase transitions, Information
Introduction
Evolutionary modeling and theory have advanced rapidly over the past century incorporating the theory of replicators [1, 2, 3, 4], sources of intragenomic variation [5, 6, 7], environmental mechanisms of genetic variation [8, 9], and socio-ecological interactions via game theory [10] and the Price equation[11].
All of these approaches rely on the concept of fitness. Typically, fitness is defined as reproductive success, estimated by ratios of offspring to parents [1, 12]. Higher relative fitness of a type results in higher prevalence in the population over time. Most studies of evolutionary dynamics adopt this framework to identify better-adapted phenotypes in natural settings in theoretical models [13, 14, 15] and laboratory studies of gene expression  [16].
However, it was pointed out that this definition of fitness is incomplete and logically circular. This is because it does not explicitly refer to environmental adaptation [17, 18]. A related issue deals with the nature of adaptive capacity, which must be a persistent property of phenotypes in variable environments and cannot depend on single events [16, 19]. Thus, any solution to this problem must include how fitness is averaged over time and environmental states. This issue also surfaces when modeling strategic interactions, such as in evolutionary game theory [20, 21].
Finally, and related to temporal and population averaging, is the question of how fitness combines across scales. It is well known in replicator models that contributions to fitness multiply [12]. This raises questions about the approximations to fitness as a linear function of traits in both modeling and empirical analysis [22, 23]. The linearity of contributions to fitness is often justified in situations of weak selection. This is commonly assumed in the analysis of kin and group selection [22, 24] and in gene expression experiments [25]. Because many results and methods are based on the linearity of fitness, it becomes important to place this assumption on a stronger basis and to consider circumstances when selection is not weak [26, 27, 28].
Other problems arise from explicit environmental or social selection, for example, in the context of evolutionary game theory. An illustration is the difficulty of deriving definite predictions from strategic interaction games for the problem of cooperation or altruism [29, 24, 30]. A related problem is the proliferation of Hamilton-type rules for similar models [31]. These complications have led to the extensive use of computer simulations with different fitness definitions [32, 33], network and group structures [34, 35], and strategy spaces [36, 37, 38, 39]. This body of work has emphasized repeated interactions [40, 41] and different types of reciprocity [31, 29] as fundamental conditions for cooperation to emerge and the need for group formation and intergroup competition [42, 22, 43]. However, these findings remain contested [24], in part because of the difficulty of analyzing the impact of assumptions and modeling choices.
Here, we show that these conceptual and practical problems share a common root in the theoretical foundations of fitness. We demonstrate how they can be addressed by re-defining fitness as an optimal statistical adaptation to a probabilistic environment. Population dynamics under selection are then identified with iterated Bayesian updates of type frequencies. This identification has been recently proposed by several authors [44, 45, 46, 47], who noted its mathematical consistency and some of its conceptual implications. We take these ideas further by showing how they lead to the interpretation of fitness as a Bayesian likelihood, which is the optimal predictor of environmental states. We also show how fitness hence constructed allows us to build models of phenotype-environment selection in simple and clear ways in a variety of complex problems. This includes the effects of probabilistic environments in task selection, evolutionary dyadic games, and selection in group-structured populations.
Phase spaces for these problems predict evolutionary transitions, expressed in the language of probability and information. Perhaps most satisfying, this approach shows how selection in evolutionary dynamics is a form of optimal learning theory in populations and how fitness maximization corresponds to the fundamental principle of maximizing information between population structure and the statistics of ecologically relevant environments.
Results
We now motivate how fitness quantifies adaptation to stochastic environments, and how this identification addresses the issues raised in the Introduction.
Population replicators and their growth dynamics
We define a structured population evolving under replicator dynamics. The population is made up of different types, $s$, such that the total population size $N=\sum_{s}N_{s}$, where $N_{s}$ is the population of each type at a given time. The temporal evolution follows the standard replicator equation for a haploid population in discrete time $\tau$,
$N_{s}(t)=w_{s}N_{s}(t-\tau),\quad{\rm so\ that}\qquad N(t)=\sum_{s}N_{s}(t)=\sum_{s}w_{s}N_{s}(t-\tau)=wN(t-\tau),$ (1)
where $w_{s}\geq 0$ is the type’s fitness. The type $s$ will grow if $w_{s}>1$ and decrease for $w_{s}<1$. This leads to the growth equation for the total population with average fitness $w=\sum_{s}w_{s}f_{s}$. Here, $f_{s}=\frac{N_{s}}{N}$ is the fraction of the population of type $s$. The total population will grow if $w>1$. The growth rate, $\gamma_{s}$, is related to fitness by exponentiation $w_{s}=e^{\gamma_{s}\tau}\simeq 1+\gamma_{s}\tau$, and, correspondingly $\gamma_{s}=\frac{1}{\tau}\ln w_{s}\simeq\frac{w_{s}-1}{\tau}$. These relationships apply in continuous time while the approximations are first-order series expansions in $\tau$, Eq. 1. In the following, we set $\tau=1$ for simplicity.
Figure 1: Scheme of evolutionary population dynamics of selection in various environments. A: The fitness as a model of adaptation becomes proportional to a likelihood of environmental states given the phenotypes. B: The iterated process over time averages over the environment statistics and selects the population types that maximize the average log-fitness as the information between the population structure and the environment.
Selection, Bayesian learning and fitness
Fitness refers to optimal adaptation to an external environment, $e$. However, Eq. 1 and the standard set up of population genetics [1] make no explicit reference to $e$ We now show how these perspectives can be reconciled .
In evolutionary dynamics, the focus of optimization is the maximization of relative fitness. The time evolution of type frequencies follows from Eq. 1 as
$f_{s}(t)=\frac{N_{s}(t)}{N(t)}=\frac{w_{s}}{w}\frac{N_{s}(t-1)}{N(t-1}=\frac{w_{s}}{w}f_{s}(t-1)\qquad\rightarrow\qquad p(H|D)=\frac{p(D|H)}{p(D)}p(H).$ (2)
A number of authors have recently noted that the replicator equation under pure selection maps exactly to Bayes relation [47, 46], written on the right side of Eq. 2. Bayes theorem gives the exact update of an initial probability that a hypothesis $h\in H$ is true, $p(h)$ (the prior), given new evidence $D$. This is written as the updated (posterior) probability $p(H|D)$. Because Bayes’ theorem is an identity that follows from the rules of probability, its procedure is the optimal way to incorporate new evidence to update the initial probability [48]. As $p(H|D)$ is more constrained than $p(H)$ the data effectively selects specific hypotheses as more likely in a process where information was learned about the hypothesis space. The key quantity in this procedure is $p(D|H)$, known as the likelihood. Although it looks like another probability, the likelihood evaluated over the data and the hypotheses is the engine of selection. It is interpreted in Bayesian estimation as a probabilistic model predicting the data given each hypothesis. The remaining factor is a normalization and does not introduce any new quantity because $p(D)=\sum_{H}p(D|H)p(H)$.
The interpretation of type frequencies as probabilities leads to the mapping of evolutionary dynamics under selection to Bayesian learning [47, 46]. This is a condition of adaptive optimality, in the sense that it gives the most predictive change in the relative population size of various types, given events in the environment [49].
This identification redefines fitness as an explicit model of environmental adaptation without additional assumptions. To see this, let us map the data, $D$, to environmental states, $D\rightarrow e$, and hypotheses to population types, $H\rightarrow s$. We then identify $f_{s}(t)=f(s|e)$, where $e$ is the state of the environment at time $t$.
$f(s|e)=\frac{p(e|s)}{{\hat{p}}(e)}f(s),\rightarrow w_{s}(e)=A(e)p(e|s),\quad w(e)=\sum_{s}w_{s}(e)f(s)=A(e)\sum_{s}p(e|s)f(s)=A(e){\hat{p}}(e).$ (3)
This leads to the identification of fitness with the Bayesian likelihood, $w_{s}(e)=A(e)p(e|s)$, up to a multiplicative constant $A(e)$. The factor of $A(e)$, is necessary because fitness, defined in Eq. 1, $w_{s}=N_{s}(t)/N_{s}(t-1)>1$, when there is growth. The factor $A$ disappears in relative fitness, but contributes to absolute fitness and to the growth rate of the entire population. Note that ${\hat{p}}(e)$ that appears in the average fitness is the estimate of the true environmental probability, $p(e)$, obtained by averaging the likelihood across types.
This achieves a definition of fitness for each type $s$ as a variable probabilistic model of the environment, addressing the circularity problem. The picture of adaptation that emerges is explicitly relational and statistical: Selection relies on each phenotype having a generative model of the environment $p(e|s)$ and it is the goodness of this model, suitably averaged, that will translate into reproductive success.
Fitness, growth rates, and information
The fitness defined in Eq. 3 is a stochastic quantity because it is evaluated at (random) states of the environment. Consequently, we must discuss how to average it over the environment and what quantities are maximized by the population dynamics of Eq. 2. The connection between fitness, probability and information comes into focus when we consider that
$\ln w_{s}(e)=\ln A(e)+\ln p(e)+\ln\frac{p(e|s)}{p(e)},$ (4)
which is the instantaneous growth rate for type $s$. The first term shows the payoffs from the environment, the second measures environmental variability. The last term measures how close the $s$-type model of the environment is to the real thing, $p(e)$. Only the last term depends on the type and contributes to relative fitness.
When the population dynamics is iterated, the types with the highest average log-fitness are positively amplified. This emerges from the temporal dynamics because
$w_{s}(E)=\prod_{t=0}^{T}A(e_{t})p(e_{t}|s),\quad\rightarrow\ {\bar{\gamma}}_{s}=\sum_{i=1}^{T}\ln A(e_{t})p(e_{t}|s)\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (5)
Here, ${\bar{\gamma}}_{s}$ is the average growth rate of type $s$ over a long time $T$ and $\ln A$ is the average of $\ln A(e)$, see Supplementary Information. The long-term temporal average becomes the average over $e$ due to the law of large numbers. Thus, the population dynamics selects the types $s$ that maximize the average growth rate, which can be interpreted in terms of information quantities [50, 51]. The relative growth rate is maximized for $s$ such that the last term (which is the relative entropy and is negative definite) vanishes. This happens when the type’s fitness $p(e|s)\rightarrow p(e)$ becomes a perfect model of the real environment. The average of the type’s growth rates over the posterior $f(s|e)$, results in the mutual information $I(E,S)$, see Supplementary Information. In this sense, selection (1-3) converges on the type distribution that maximizes the information between population structure and the statistics of the environment.
Two additional consequences follow from the definition of fitness as a Bayesian likelihood. First, because fitness is a probability, statistically independent contributions combine as products. This means that fitness will only rarely be a linear combination of factors as is often assumed. Second, independent contributions do add up in the log-fitness, as they do in all information quantities. This makes the average log-fitness (or growth rate) the natural quantity for the analysis of complex evolutionary problems, even away from weak selection.
Fitness in probabilistic environments
We now illustrate the connections between population dynamics, Bayesian selection, and information through the analysis of several canonical models. We show that this approach creates useful analytic tools for evolutionary game theory and the derivation of Hamilton-type rules for the emergence of altruistic interactions in structured populations.
Task switching
In the problem of task switching, a population of agents is distributed to a set of discrete tasks, $s$, each associated with differential payoffs dependent on environmental states, $e$. Optimal evolutionary dynamics align these tasks with the environment to maximize rewards. For simplicity, we consider two tasks, denoted $s=0,1$, in a binary environment, $e=0,1$. The standard setup prescribes rewards, $Y_{s}(e)$, to each of these state combinations, Table 1A-B.
To understand the dynamics of selection, note that it pays off to align the task with the environment $s=e$, regardless of the specific state, $e=0,1$ because $Y_{s=e}=Y_{+}>Y_{s\neq e}=Y_{-}$. The question becomes interesting when the environment is only known in probability, $p(e)$. Then, starting with some probability of tasks $f(s)$, the population is selected over time, updating the conditional probability $f(s|e)$ as instances of the environment are experienced.
The identification of population dynamics of this type with iterated Bayesian selection requires mapping the rewards $Y_{s}(e)$ to the likelihood function, $p(e|s)$. Because the likelihood is a conditional probability model, it is positive and normalized, meaning that any mapping to payoffs must satisfy such requirements. We want mappings that preserve the relative strength of the payoffs. This is achieved by writing the fitness, $w_{s}(e)=Ap(e|s)=Y_{s}(e)$. This requires that $Y_{s}(e)>0$ as assumed in Table 1 and that the factor of $A(e)$ is defined to ensure that none of the probabilities exceeds one and that the conditional probability is normalized. In general, the factor of $A$ is a function of the state of $e$, but since the example of Table 1 is symmetric, $A=5/2$ and $p(e=0|s=e)=p(e=1|s=e)=\frac{4}{5}$ and $p(e=0|s\neq e)=p(e=1|s\neq 0)=\frac{1}{5}$. We can write the fitness as
$w_{s}(e)=Y_{+}^{(1-s)(1-e)+se}Y_{-}^{(1-s)e+s(1-e)},\qquad\ln w_{s}=\ln Y_{+}+\ln\frac{Y_{+}}{Y_{-}}\left(2se-s-e\right).$ (6)
Note that we can now obtain the probability of $s$ that will maximize fitness over a long (stationary) history of the environment by taking expected values over $e$ and $s$. We define $f(s=1)=p$ and $p(e=1)=q$, and $f_{1|1}=f(s=1|e=1)$ and $f_{1|0}=f(s=1|e=0)$. The log-fitness is maximized when the quantity $(2f_{1|1}-1)q-p$ is maximal. This requires $f_{1|1}=1$, which implies $p=f_{1|1}q+f_{1|0}(1-q)=q$, which means that the population distribution of tasks is equal to the environmental probability.
We can also write the dynamics of selection by iterating the likelihood, given an environmental history and an initial distribution of traits $s$ in the population, see Supplementary information
$f(s|e)\sim\Pi_{t=1}^{T}p(e_{t}|s_{t})f(s)=Y_{+}^{N_{e=s}}Y_{-}^{T-N_{e=s}}f(s),$ (7)
where $N_{e=s}$ is the number of times that $s=e$. For large times, the law of large numbers tells us that $N_{e=s}\rightarrow Tp(e,s)=Tp(s|e)p(e)$. The likelihood is maximized by taking $p(s=e|e)\rightarrow 1$, which means that $f(s)\rightarrow p(e)$, as derived above.
We conclude that the state of the population that maximizes the (Bayesian) dynamics of selection exactly mirrors that of the environment. The state of $s$ selected at long times also maximizes the information that the population has on the environment. This follows from $I(S,E)=H(S)-H(S|E)=H(E)$, meaning that the variation of the environment is now fully explained by population statistics. In this sense, evolutionary dynamics selects the state that maximizes information transmission between the environment and the population.
Evolutionary games
Evolutionary game theory provides a general framework for the study of coevolutionary dynamics resulting from social interactions. The population now becomes an environment from which each agent derives its fitness. This contrasts with the task selection problem, where environmental dynamics are stationary and therefore independent of agent actions. The simplest situation is a mixture of two games: One game when the state of the environment $e=0$, and another when $e=1$. The total payoff over time will then be the result of agents recognizing which environmental situation they are in and adjusting their interactions accordingly. This also generalizes the task-switching problem to what we may call a game-switching situation.
We first demonstrate how the statistical treatment of fitness simplifies the analyses of dyadic games. Consider first the prisoners’ dilemma (PD), used for decades to analyze the evolution of cooperation [29, 20]. Table 1C-D shows the game’s payoff structure, where $s=0$ corresponds to defection and $s=1$ to cooperation. We consider two players $s_{1},s_{2}$ each with probability cooperation $p_{1},p_{2}$, respectively. Payoffs are chosen such that $W>R>L>P>0$. We write $s_{1}$’s fitness as
$w_{s_{1}}(s_{2})=AR^{s_{1}s_{2}}W^{(1-s_{1})s_{2}}P^{s_{2}(1-s_{1})}L^{(1-s_{1})(1-s_{2})}.$ (8)
Note that there is no unique prescription for the population dynamics of games [52]. Eq. 8 gives stochastic pay-offs at each interaction rather than on average, which would requires additional assumptions [52, 53]. This leads to
$\ln w_{s_{1}}=a-cs_{1}+bs_{2}+ds_{1}s_{2},\qquad a=\ln AL,\ c=\ln\frac{L}{P},\ b=\ln\frac{W}{L},\ d=\ln\frac{RL}{PW}.$ (9)
This form of log-fitness coincides with the typical parameterization of fitness as a linear function of costs and benefits of cooperation in the classical problem of altruism. Specifically, there is a fitness cost $c>0$ of cooperating when $s_{1}=1$, an independent benefit $b>0$ from the cooperation of others, $s_{2}=1$. There is also an extra benefit (if $d>0$) when both agents cooperate, which was considered by Queller [23, 22]. Note also that $d=c-\delta<c$, with $\delta=\ln W/R>0$. We see that, in a single-shot game, $s_{1}=0$ (defection) results in higher log-fitness than cooperation, even when the other player cooperates ($s_{2}=1$). This is the familiar Nash equilibrium, because the log-fitness of $s_{2}$ is identical with labels interchanged.
A Fitness, $w_{s}(e)$ $e=0$ $e=1$ $s=0$ $Ap(0|0)$ $Ap(1|0)$ $s=1$ $Ap(0|1)$ $Ap(1|1)$ B Returns, $Y_{s}(e)$ $e=0$ $e=1$ $s=0$ $Y_{+}$ $Y_{-}$ $s=1$ $Y_{-}$ $Y_{+}$ C $Y_{s_{1},s_{2},e=0}$ $s_{2}=0$ $s_{2}=1$ $s_{1}=0$ $L,L$ $W,P$ $s_{1}=1$ $P,W$ $R,R$ D $Y_{s_{1},s_{2},e=1}$ $s_{2}=0$ $s_{2}=1$ $s_{1}=0$ $L,L$ $P,W$ $s_{1}=1$ $W,P$ $R,R$
E       Relation Cooperation Condition Parameters Statistical independence (SI) never Well-mixed population (WM) $r_{\rm m}\leavevmode\nobreak\ b+f_{\rm m}\leavevmode\nobreak\ d>c$ $r_{\rm m}=1$, $f_{\rm m}=p$ Conditional dependence (CD) $r_{\rm d}\leavevmode\nobreak\ b+f_{\rm d}\leavevmode\nobreak\ d>c$ $r_{\rm d}=f_{1|1}-f_{1|0}$, $f_{\rm d}=f_{1|1}$ Mixed games $r_{\rm d}\leavevmode\nobreak\ b(q)+f_{\rm d}\leavevmode\nobreak\ d>c(q)$ $q=p(e=1)$ Selection in groups $\frac{k_{g}}{K}\left(r_{g}\leavevmode\nobreak\ b+f_{g}\leavevmode\nobreak\ d\right)+dp_{\bar{g}}>c$ $r_{g}=f_{1g|1g}-f_{1g|0g}$, $f_{g}=f_{1g|1g}-p_{\bar{g}}$
Table 1: Combined payoff tables for the task switching problem and 2-agent games. A: Payoffs for task alignment with the environment, where $Y_{+}=2,Y_{-}=1/2$. B: Fitness mapping to likelihood. C: Payoffs for the prisoners’ dilemma ($e=0$) with $W>R>L>P$. D: Payoffs for the cooperation game ($e=1$) with $W^{\prime}=P,\ P^{\prime}=W$, inverting the prisoner’s dilemma cross strategy payoffs. E: Summary of conditions for the emergence of cooperation.
By contrast, the goal of evolutionary game theory is the identification of selected strategies over long histories of interaction, which is analyzed via averaging Eq. 9. There are three distinct situations with growing levels of complexity. First, when the two agents act independently. Second, when they are part of a fully mixing population. And third, when their phenotypes become (statistically) dependent. Each of these general situations is associated with different population outcomes.
When the agents are statistically independent with cooperation probabilities $p_{1},p_{2}$, the average log-fitness is
$\ln w_{s_{1}}=a-cp_{1}+bp_{2}+dp_{1}p_{2}=a-\left[c(1-p_{2})+\delta p_{2}\right]p_{1}+bp_{2}.$ (10)
The quantity in the square bracket is always positive, so $p_{1}=0$ maximizes the fitness independently of $p_{2}$. This means that the agent should always defect, as expected. Here, statistical independence means intuitively that the other agent does not react, even when it is exploited. In this sense, it must correspond to the single shot game and has the same optimal solution. Note also that, because of statistical independence, there can be no information between the agents.
Now consider a seemingly small but important change in these assumptions. We consider $p_{1}=p_{2}=p$, which is frequently assumed in population genetics. It is dubbed random mating (in diploid populations) or a fully mixing population [1]. This results in
$\ln w_{s_{1}}=a-(c-b)p+dp^{2},$ (11)
The polynomial form of the log-fitness in $p$ is familiar and simple to analyze. A distinction emerges depending on the relative magnitude of $b$ and $c$, Fig. 2A. When $b>c$ (or $WP>L^{2}$) the situation is simple because the log-fitness is always a growing function of $p$, with a single maximum at $p=1$ corresponding to full cooperation. Whatever the initial state of the population, cooperation will always emerge as the dominant strategy and will eventually fixate. Relative to the statistically independent case, cooperation now emerges from the fact that the population is effectively playing itself, so that increases in mutual exploitation lead to overall lower fitness. The situation is more interesting when $c>b$, because this leads to a fitness minimum at $p_{*}=\frac{c-b}{d}<1$. The existence of this fitness minimum creates a local fitness maximum at $p=0$ (mutual defection), which is locally stable. In these circumstances, sufficient cooperation $p>p_{*}$ must be present for the population to converge to the global fitness maximum of full cooperation ($p=1$). This corresponds to a simple condition for cooperation to emerge, related to Hamilton’s rule [54]. From Eq. 11, we identify benefits of cooperation as $B=b+dp$ and costs with $C=c$. Then $B>C$ is a condition for full cooperation to emerge, which is equivalent to $p>p_{*}$. The population dynamics around $p=p_{*}$ is characterized by a first-order transition – a "tipping point"– because there is a valley of fitness to be crossed to reach the global maximum, starting with generalized defection. This can be done locally, over space or on networks, as many authors have noted based on numerical experiments [27, 55]. We will return to this point below.
Third, the situation is both more interesting and more complex when the agents’ actions become statistically dependent. In this case, there are three parameters at play in the joint distribution: four joint probabilities minus the normalization. We can parameterize this space in analogy to task selection as $p_{2}=f_{1|0}(1-p_{1})+f_{1|1}p_{1}$ and $f(s_{1}=1,s_{2}=1)=f_{1|1}p_{1}$, with the shorthand $f_{1|1}=f(s_{2}=1|s_{1}=1)$, $f_{1|0}=f(s_{2}=1|s_{1}=0)$. Note that $r=f_{1|1}-f_{1|0}$ is the regression of $p_{2}$ on $p_{1}$, which vanishes when the two agents are statistically independent. Thus, $r$ is a measure of how (co-)related the two states are. The log-fitness becomes
$\ln w_{s_{1}}=a+bf_{1|0}+[rb+df_{1|1}-c]p_{1}.$ (12)
Because this is linear in $p_{1}$ the nature of the maximum depends only on the sign of the term in the square brackets. This is a second version of Hamilton’s rule closer to the way Queller [22] expressed it. If we identify the benefits of interaction $B=b$, and costs $C=c$ we now obtain $rB>C$ as the condition for cooperation to emerge. Adding the $d$ term facilitates the emergence of cooperation via what Queller called a synergism coefficient, as $rB+df_{1|1}>c$. This transition has a geometric interpretation in the reciprocity-exploitation ($f_{1|1},f_{1|0}$) plane. We can write the cooperation condition as a line, $y=f_{1|1}$, $x=f_{1|0}$, $y=y_{0}+\beta x$, with $y_{0}=\frac{c}{b+d}$, $\beta=\frac{b}{b+d}$ , as shown in Figure 2C. Cooperation, $p_{1}\rightarrow 1$, emerges for values above this line. To be above the line requires controlling exploitation, optimized when $f_{1|0}=f(s_{2}=1|s_{1}=0)\rightarrow 0$ (which implies that punishment or retaliation, $f(s_{2}=0|s_{1}=0)\rightarrow 1$. It also requires reciprocity meaning that $f_{1|1}=f(s_{2}=1|s_{1}=1)\rightarrow 1$ ( so that one becomes safe from exploitation, $f(s_{2}=0|s_{1}=1)\rightarrow 0$). Note that the first of these conditions may take the population towards mutual defection  [56], which is suboptimal. The second is therefore necessary to reach the highest fitness globally. Below the line, cooperation cannot emerge. These outcomes agree with established results on the effects of punishment and rewards on cooperation [57, 58]. That is, both punishment and reciprocity promote the establishment of cooperation but alone punishment does not suffice [59]. The information in the population is maximal as agents coordinate but the amount of information at stake converges to zero at full cooperation because then there is no uncertainty left to explain.
Other 2x2 games can be analyzed in the same way, see Supplementary Information. Phase transitions are expected as the polynomial terms in the log-fitness change sign driven by external parameters such as characteristics of the environment, as we show next.
Figure 2: Phase diagram for the Prisoner’s Dilemma and the environmentally mixed game. A: Log-fitness versus cooperation probability $p$ for the well-mixed population. Note the fitness valley for $b<c$. Conditional dependent case conditions for cooperation. B: Probability of reciprocity $f_{1|1}$ versus environment, $q$. C: Probability of reciprocity versus exploitation $f_{1|0}$, for various $q$. Cooperation emerges in each case for parameters above the lines. See also Fig.S1
Mixed games in variable environments
Greater realism and more possibilities result from combining several interaction strategies presented by the environment. Here we illustrate this situation via the mixture of two games, PD for $e=0$, and "turn the other cheek" (ToC), for $e=1$. ToC is a game of virtue, where it is shameful (costly) to exploit a cooperator and the greatest reward is obtained by cooperating with a defector. It is the game obtained by inverting PD’s off-diagonal elements, Table 1C-D.
Selected strategies depend on the relative probabilities of the two games set by the environment $q=p(e=0)$. The log-fitness becomes
$\ln w_{s_{1}}=a-c(e)s_{1}+b(e)s_{2}+ds_{1}s_{2},\quad{\rm with}\quad c(e)=c(1-e)-be,\ b(e)=b(1-e)-ce.$ (13)
Costs and benefits swap between the PD and the ToC, for which $\ln w_{s_{1}}=a+bs_{1}-cs_{2}+ds_{1}s_{2}$. If the two agents act independently, then the agent should always cooperate in the ToC because both $b,d>0$. If the two agents act independently but share the same probability $p_{1}=p_{2}=p$, the log-fitness is independent of $e$ because there is a symmetry between the two agents so that the benefits and costs of cooperation are fully socialized. Thus, in the fully mixed population, changes in the environment have no effect on the average population dynamics and the phase diagram is the same as in Fig.2B.
When the two agents coordinate, the situation becomes more interesting. If the agents can tell which game they are playing at each time, their probability of cooperation will be the selected strategy for each game weighed by their relative probability $q$. The other situation we consider is when agents act conditionally on each other independently of the state of the environment. Then we obtain the same condition for cooperation to emerge as in PD, but with costs and benefits that depend on the environment as in Eq. 13. Fig. 2B-C shows that when $q=0$, we obtain the PD condition (black line) in Fig. 2. As $q$ increases, both the ordinate at the origin and the slope decrease. At $q=1/2$ the slope is zero. Eventually, at $q=2/3$, ToC is sufficiently dominant and cooperation becomes unconditional.
This analysis is general and can be obtained for any combination of games leading to complex and interesting phase spaces, Fig. 2. As a result, we obtain transitions between different (optimal) population states as functions of environment probabilities, which play the role of control parameters, analogous to temperature in physical systems. We have therefore shown how mixing the prisoner’s dilemma with a cooperative game increases the selection for cooperation in non-trivial ways. This is likely a more realistic situation, as organisms and especially people, have a diversity of modes of interaction some competitive and some cooperative. Thus, the presence of explicitly cooperative channels some of the time can help fix unstable cooperation all of the time.
Selection in groups
The probabilistic interpretation of fitness readily extends to structured populations. Within a population, group identities specify additional layers of organization that condition phenotypes and therefore may contribute to individual selection.
We write the fitness of individual $i$ in group $g$ as $w_{gi}(g^{\prime},j)=Ap(\{s_{g^{\prime}j}\}|s_{gi})$, where $\{s_{g^{\prime}j}\}$ denotes the joint state of all neighbors $j,g^{\prime}$. Typically, these neighbors are not statistically independent from each other leading to synergistic interactions[23, 60]. This is important for finding clusters of states (genes, phenotypes) that collectively encapsulate adaptive functions [61].
However, the most common situation arises when interactions are dyadic (between pairs, only), so that the game occurs on a simple graph [62, 63, 35, 64]. In the log-fitness, this approximation is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that
$w_{gi}=\Pi_{j,g^{\prime}}Ap(s_{g^{\prime}j}|s_{gi}),\rightarrow\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (14)
With these assumptions, we can write the log-fitness to be identical to Eq. 10, where the second player is replaced with the average over all connections, $s_{2}=1/K\sum_{g^{\prime},j}{s_{g^{\prime}j}}$. All our previous results still apply but require attention to group dependencies (see Supplementary Information for details).
Figure 3: Selection in structured populations depends on group sizes and treatment of out-groups. A: Phase diagram as a function of cooperation from outer groups $p_{\hat{g}}$ and fraction of connections within in-group, $k_{g}/K$, for various amount of in-group relateness $r$. B: Phase diagram as a function of cooperation probability in out-groups and in group reciprocity, for different fractions if in-group connectivity, $k=k_{g}/K$. Cooperation emerges for conditions above each line. C: Treating out-groups as independent promotes defection across groups and in the general population. D: Cooperation can be expanded as interactions become cooperative across group labels.
In particular, the agent should still defect unconditionally when it is independent from the other agents. When their probabilities coincide, in the well-mixed population $p_{gi}=p_{2}=p$, Eq. 11 still applies and a phase transition occurs at $p=p_{*}$, if $b<c$. When the states are not independent, the log-fitness is written compactly in terms of contributions from each interaction, weighed by the size of the population $1/K$,
$\frac{1}{K}\ln w_{gi}=a+bf_{1|0g}+[rb+df_{1|1g}-c]p_{gi}.$ (15)
This equation is similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$, where the additional $g^{\prime},g$ labels denote conditional dependence across groups. The absence of the label means average across groups. The interesting possibility is that the agents’ behaviors are now different within and across groups in the sense that reciprocity $f_{1g^{\prime}|1g}<f_{1g|1g}$, and exploitation $f_{1g^{\prime}|0g}>f_{1g|0g}$, for $g^{\prime}\neq g$. Treating out group interactions as statistically independent leads to
$\ln w_{gi}=a+\underbrace{k_{g}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}+\underbrace{\sum_{g^{\prime}\neq g}k_{g^{\prime}}\big{[}\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\big{]}}_{\rm out-groups}.$ (16)
This presents an interesting conundrum: The in-group contribution drives cooperation, but the out-group selects for defection. In general, the out-group is larger so cooperation may be corrupted by the temptation to exploit out-group members, or simply by the failure to coordinate phenotypes.
This collapse of cooperation under external pressure can be averted in two ways. First, the agent may cease to interact with out-group members so that $k_{g^{\prime}\neq g}=0$, but this foregoes fitness benefits from the possible altruism of a large population. Fig.3A shows when these benefits promote cooperation, solving $rB>C$ for $k_{g}/K$ as a function of $p_{\bar{g}}$. As before, the contours define a boundary, above which more cooperation increases fitness. We also plot this solution in Fig3B for $f_{1g|1g}$ and $p_{\bar{g}}$, which can be compared to Fig.2B when $q=0$.
Second, cooperation or defection could conditionally depend on the group identity of the interaction partner. Then the in-group and out-group probabilities of cooperation $p_{gi|g^{\prime}}$ adapt differentially so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$. This may still mean that the average cooperative behavior $p_{gi}=\frac{k_{g}}{K}p(gi|g)+\frac{K-k_{g}}{K}p(gi|g^{\prime}\neq g)$ is small if most interactions are external, as illustrated in Fig.3A, but cooperation will now be stabilized within each group. We note in Fig.3B that this need not be the final situation. Coordination and cooperation can continue to spread across groups selectively, as larger sets of group markers form new collective identities associated with in-group behavior through the mechanisms described in Table 1D. In this way, pedigree relations may be the most natural basis for in-group formation [22], but mutualistic strategic interactions can also spread if they become associated with recognizable characteristics and beneficial phenotypic behaviors.
Discussion
We have demonstrated how to resolve several conceptual and practical problems in evolutionary theory by modeling fitness as an explicit process of adaptation to statistical environments. This result arises from interpreting population type frequencies as probabilities and assuming that adaptation is an optimal encoding of environmental evidence through Bayesian inference–a standard mapping in evolutionary dynamics [47, 46].
Writing fitness as a Bayesian likelihood clarifies the construction of mathematical models of adaptation in terms of conditional probability functions. This definition shows that fitness combines across scales following the rules of probabilities. This allows us to identify statistically independent contributions to fitness (from multiple alleles, population structures or other factors) as additive in log-fitness, which is not a property of fitness itself except in the limit of weak selection. Then, the long-term population dynamics of selection maximize average log-fitness, which is the maximal mutual information between the environment and population structure. In this sense, the guiding principle of natural selection –fitness maximization– corresponds to choosing the population distribution across types that maximizes the information between the environment and its structure [49]. This occurs gradually, as the result of long-term adaptive population dynamics under selection.
We illustrated the practical advantages of this approach to fitness in well-known evolutionary theory problems. This allowed us to analyze the task selection problem, evolutionary games when interactions change as a function of the environment, and selection in groups. We showed that conditional probabilities define key mechanisms in evolutionary games ranging from population assortment to reputation and punishment [39] thereby expressing general biological and social processes in terms of statistical quantities.
The approach and results of this paper are part of a broad effort in the literature to make information more central in biology and evolution [65, 49, 19, 18, 45, 47, 14]. To this end, many authors have identified information theory as a natural language for biological phenomena ranging from the evolution of genome organization[65, 66], the emergence of structure in ecological networks[67], and fitness as it relates to environmental stochasticity [12, 14]. By formalizing the environment as a learnable statistical model, our approach replaces the circular notion of fitness as reproductive success by a well-defined statistical quantity that is optimized through recursive inference and characterized by information. In introducing this mechanism, we open up several directions for future research.
First, defining fitness as a likelihood can guide both field observations and laboratory studies of evolutionary dynamics. This approach emphasizes identifying environmental states $e$ that yield fitness payoffs via measurement of $A$, and their statistical relations to specific phenotypes, $s$. This approach reduces (or compresses, in probability) the dimensionality of large spaces of environmental states to those with fitness consequences when conditioned on given phenotypes. Statistically independent factors can be identified as additive (in multiple variable linear regression) in the log-fitness (average growth rates). Practical applications in empirical settings also require identifying sources of variation not associated with selection. In standard evolutionary theory, mutations deviate population dynamics from pure Bayesian updating. Although fitness is still updated through the likelihood function, the dynamics now reflect a mutation-selection balance. Non-selective forces such as genetic drift and gene flow prevent the complete optimization of information transfer between the population and the environment, leading to incomplete adaptation and fixation of dominant types. These factors introduce parameters, such as the magnitude of fluctuations, to the analysis which will typically increase the evolvability of populations and enrich phase-space possibilities for evolutionary transitions.
Many of these possibilities will lead to increased group heterogeneity. Our analysis of dyadic interactions shows that clusters of related phenotypes can promote cooperative behavior, introducing an information signal that can drive selection as argued in the general problem of major evolutionary transitions [68]. This approach generalizes beyond pairwise interactions, such as complex divisions of labor. This allows collective interactions to be analyzed via the structure of conditional (in)dependence of states, as established in probabilistic graphical models [69]. These results strengthen the intersection of evolutionary biology and statistical learning theory by integrating evolutionary dynamics with causal inference, information theory, and statistical learning, thereby paving the way for interdisciplinary applications in artificial intelligence.
Methods
Python code for generating the phase diagrams in Figures 2-3 is available online at https://zenodo.org/records/15003106.
Acknowledgements
We thank Stefano Allesina, Linnéa Gyllingberg, David Krakauer, Michael Lachner, and Ruy Ribeiro for useful discussions and Duncan Rocha for feedback on the figures. This work is supported by the Conard Initiative on the Foundations of Complexity (to L.M.A.B), the Institute for New Economic Thinking at the Oxford Martin School (to J.T.K), and the National Science Foundation Graduate Research Fellowship Program (Grant No. DGE 2140001 to B.J.G.)
Author contributions statement
L.M.A.B. conceived the research topic and developed the main theory. L.M.A.B. and J.T.K. developed the experiments and conducted the analysis. L.M.A.B. drafted the manuscript. B.J.G. conducted the literature review and J.T.K. produced the figures. All authors reviewed the manuscript.
Supplementary Information
Average growth rates and Information
Here, we give additional details about the connection between population growth rates and information. In Eq. 4 we wrote the log-fitness, and equivalently the instantaneous growth rate, as
$\gamma_{s}\tau=\ln w_{s}(e)=\ln A(e)+\ln p(e|s)=\ln A(e)+\ln p(e)+\ln\frac{p(e|s)}{p(e)}.$ (S.1)
Below, we take $\tau=1$ for simplicity. The population dynamics expose these instantaneous quantities to a sequence of states of the environment, resulting in averaging over $p(e)$. This is due to the law of large numbers
${\bar{\gamma}}_{s}=\sum_{t=1}^{T}\ln\big{[}A(e_{t})p(e_{t}|s)\big{]}\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (S.2)
which is also the average log-fitness. We can write these averages in terms of information quantities, leading to
${\bar{\gamma}}_{s}=\ln A-H(E)-D\big{(}p(E)||p(E|s)\big{)}$ (S.3)
Here $\ln A=\sum_{e}p(e)\ln A(e)$, $H(E)=-\sum_{e}p(e)\ln p(e)$ is the entropy of the environment, and the $D\big{(}p(E)||p(E|s)\big{)}\geq 0$ is the relative entropy or Kullback-Leibler divergence, which is positive definite. It follows that the average growth rate is maximized when $D$ is minimized at $p(e|s)=p(e)$. This means that type $s$ has a perfect statistical model of its environment. The other two terms are also interesting but do not affect the relative fitness because they are independent of $s$. They show that the magnitude of the payoffs $\ln A>H(E)$ for there to be growth, and that environments of decreasing complexity, $H(E)\rightarrow 0$, are more easily predicted and yield higher average rewards. These results are known in statistical models of gambling and bet-hedging.
Finally, we note that when we average the $s$-type growth rates over the new state of the population (posterior) we obtain
$\sum_{s,e}f(s|e)p(e)\ln w_{s}=\ln A-H(E)+\sum_{s,e}f(s|e)p(e)\ln\frac{p(e|s)}{p(e)}=\ln A-H(E)+I(E,S),$ (S.4)
where $I(E,S)$ is the (mutual) information between the population type frequencies and their environment. This shows that maximal population growth across types corresponds to the population structure that maximizes the mutual information $I(E,S)$, as indicated in Figure 1.
Analysis of the Bayesian Task Switching Problem
Task-switching can be easily analyzed by inspection. In this problem, the probability of the environment $p(e)$ is given, along with the likelihood (fitness) which maps to the reward matrix as explained in the main text. The state of the population in each task, $s$, is then the dynamical variable. We start with some arbitrary initial population state $f(s)$, and iterate to create a posterior $f(s|e)$. Iteration creates a posterior with a longer and longer history of exposure (selection) from the environment, starting with $f(s)$. Because the problem is symmetric in the two task states, neither is preferred and it only matters if the task is aligned with the state of the environment $s=e$ or not. We can see this by explicit inspection
$f(s=1|e)=\frac{p(e|s=1)}{{\hat{p}}(e)}f(s=1)=\frac{p(e|s=1)}{p(e|s=1)f(s=1)+p(e|s=0)f(s=0)}f(s=1).$ (S.5)
Considering the two possible states of the environment and using Table 1 leads to
$f(s=1|e=1)=\frac{p(e=1|s=1)}{p(e=1|s=1)f(s=1)+p(e=1|s=0)f(s=0)}f(s=1)=\frac{4}{1+3f(s=1)}f(s=1).$ (S.6)
Because the relative fitness $\frac{4}{1+3f(s=1)}\geq 1$, the conditional posteriors will converge to unity. Conversely, the unaligned probability
$f(s=1|e=0)=\frac{p(e=0|s=1)}{p(e=0|s=1)f(s=1)+p(e=0|s=0)f(s=0)}f(s=1)=\frac{1}{4-3f(s=1)}f(s=1).$ (S.7)
Now, the relative fitness $\frac{1}{4-3f(s=1)}\leq 1$, so that the posterior probability always reduces. By symmetry, the results hold when $s=0$. Thus, the long-term limit of the dynamics is $f(s=e|e)=1,\ f(s\neq e|e)=0$. This result enables us to determine the equilibrium state of the marginal, $f(s)$
$f(s)=f(s=e|e)p(e)+f(s\neq e|e)p(e)=p(e).$ (S.8)
which means that the task is fully aligned with the environment, in probability. This result also follows directly from the interpretation of the average growth rate as information and its maximization, as discussed in the main text.
Fitness of a general dyadic game
Here we show that any 2x2 game can be written in the form of Table 1. Consider the payoff matrix $Y_{ij}$, where the first index corresponds to the focal agent and the second to the opponent. Each player has two states $0,1$. Then
$w_{s_{1}}(s_{2})=AY_{00}^{(1-s_{1})(1-s_{2})}Y_{01}^{(1-s_{1})s_{2}}Y_{10}^{s_{1}(1-s_{2})}Y_{11}^{s_{1}s_{2}},$ (S.9)
where $w_{s_{1}}(s_{2})$ represents the fitness of a player in state $s_{1}$ when the opponent is in state $s_{2}$. Thus, the log-fitness can be written as
$\ln w_{s_{1}}(s_{2})=\ln Y_{00}+\ln\frac{Y_{10}}{Y_{00}}s_{1}+\ln\frac{Y_{01}}{Y_{00}}s_{2}+\ln\frac{Y_{00}Y_{11}}{Y_{01}Y_{10}}s_{1}s_{2}.$ (S.10)
This quantity is always a polynomial in the stochastic variables $s_{1},s_{2}$. Its polynomial coefficients will be positive or negative depending on the specific game structure, leading to different phase diagrams as shown in the main text. Averaging the log-fitness is simple. It consists of taking expectation values of each variable and their products.
Analysis of selection in groups
Here, we provide additional mathematical details about group structure and resulting patterns of selection. First, we note that the interpretation of fitness as a likelihood function extends readily to groups. Groups can be considered coordinates of an additional layer, or level, which can modulate local parameters, as is common in (neural) networks. For example, a group may have greater reciprocity with its own elements than with other groups.
We write the fitness of individual $i$ in group $g$, in interaction with other individuals $j$ in groups indexed by $g^{\prime}$ as
$w_{g,i}=Ap(\{s_{g^{\prime}j}\}|s_{gi}),$ (S.11)
where $\{s_{g^{\prime}j}\}$ denotes the set of neighboring individuals indexed by $j,g^{\prime}$. In general, these neighbors are not statistically independent from each other, leading to a property known as synergy [60]. This is important for finding clusters of states (genes, phenotypes) that collectively encode adaptive functions. This phenomenon arises, for example, when there are divisions of labor and coordination among states [61] toward collective goals.
However, the situation most commonly analyzed is when the neighbor interactions are described by a common graph (or network) reducing all interactions to dyadic ones. This is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that
$w_{gi}=A\prod_{g^{\prime},j}p(s_{g^{\prime}j}|s_{gi}),\rightarrow\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (S.12)
Thus, conditionally dependent, higher order, clusters of $s_{g^{\prime}j}$ will not be additive in the log-fitness. Generally, additivity provides a statistical test for the existence of higher level functions and structures. An intermediate situation is when the group structure is expressed as conditional independence across groups. In this case, we have
$w_{g,i}=Ap(\{s_{g^{\prime}j}\}|s_{gi})=A\prod_{g^{\prime}}p(g^{\prime}|s_{gi}),\ {\rm with}\ p(g^{\prime}|s_{gi})=\prod_{j\in g^{\prime}}p(\{s_{g^{\prime}j}\}|s_{gi})\rightarrow\ln w_{gi}=\ln A+\sum_{g^{\prime}}\ln w_{gi}(g^{\prime}),$ (S.13)
with $\ln w_{gi}(g^{\prime})=\sum_{j\in g^{\prime}}p(\{s_{g^{\prime}j}\}|s_{gi})$ the group log-fitness, obtained by averaging the fitness over interactions with all members of group $g^{\prime}$.
If we take $\ln p(s_{g^{\prime}j}|s_{gi})=a-cs_{gi}+bs_{g^{\prime}j}+ds_{gi}s_{g^{\prime}j}$, generalizing the 2-agent expression in the main text, we can write the log-fitness in terms of the sums of neighbors to our focal node. In general, these sums depend on the group structure and specifically on how interactions vary within and across groups. The relative size of these groups then becomes an important factor. We take the size of group $g$ to be $k_{g}$, and the total number of connections (neighbors) to be $K=\sum_{g}k_{g}=N-1$, where $N$ is the total population size across all groups $N=\sum_{g}N_{g}$, and $-1$ subtracts self-interactions. Many rich and complex situations can be explored as these possibilities of size and interaction structure are varied. Within our assumptions, we can then write the log-fitness as
$\ln w_{g,i}=aK-cKs_{gi}+b\sum_{g^{\prime},j}s_{g^{\prime}j}+ds_{gi}\sum_{g^{\prime},j}{s_{g^{\prime}j}}.$ (S.14)
We see that Eq. S.14 is identical to Eq. 10, as
$\frac{1}{K}\ln w_{gi}=a-cp_{gi}+bp_{2}+dp_{gi}p_{2}$ (S.15)
with the identification $p_{gi}=p(s_{gi}=1)$ and
$p_{2}=\frac{1}{K}\sum_{g^{\prime}j}p(s_{g^{\prime}j}=1)=\frac{1}{K}\sum_{g^{\prime}}k_{g^{\prime}}p_{g^{\prime}},\qquad p_{g^{\prime}}=\frac{1}{k_{g^{\prime}}}\sum_{j\in g^{\prime}}p(s_{g^{\prime}j}=1).$ (S.16)
The conditions for the emergence of cooperation can now be assessed within and across groups using the results in the main text. First, the agent should defect unconditionally when the other agents are independent of itself. In the well-mixed population, when all probabilities of cooperation are the same, $p_{gi}=p_{2}=p$, Eq. 11 still applies and a phase transition occurs at $p=p_{*}$, if $b<c$.
The situation is more interesting when states are not statistically independent and reciprocity and exploitation become group-specific. We write the generalization of Eq. 12 with
$p_{2}=\frac{1}{K}\sum_{g^{\prime}j}\left[f(s_{g^{\prime}j}=1|s_{gi}=0)+r_{g^{\prime}j}p_{gi}\right],\quad r_{g^{\prime}j}=f(s_{g^{\prime}j}=1|s_{gi}=1)-f(s_{g^{\prime}j}=1|s_{gi}=0).$ (S.17)
Note that, in the statistical independence limit $r_{g^{\prime}j}\rightarrow 0$ and the probabilities become unconditional. By averaging these conditional probabilities over the members of each out-group
$f_{1g^{\prime}|1g}=\frac{1}{k_{g^{\prime}}}\sum_{j\in g^{\prime}}f(s_{g^{\prime}j}=1|s_{gi}=1),\ f_{1g^{\prime}|0g}=\frac{1}{k_{g^{\prime}}}\sum_{j\in g^{\prime}}f(s_{g^{\prime}j}=1|s_{gi}=0),$ (S.18)
we can write the fitness in terms of the reciprocity with each group, with $r_{g^{\prime}}=f_{1g^{\prime}|1g}-f_{1g^{\prime}|0g}$, as
$p_{2}=\frac{1}{K}\sum_{g^{\prime}}k_{g^{\prime}}\left[f_{1g^{\prime}|0g}+r_{g^{\prime}}p_{gi}\right].$ (S.19)
We can repeat this again, by averaging across all out-groups
$f_{1|1g}=\sum_{g^{\prime}}\frac{k_{g^{\prime}}}{K}f_{1g^{\prime}|1g},\quad f_{1|0g}=\sum_{g^{\prime}}\frac{k_{g^{\prime}}}{K}f_{1g^{\prime}|0g},$ (S.20)
with the reciprocity defined over the average across all out-groups, $r=f_{1|1g}-f_{1|0g}$. Finally, we can write the fitness simply in terms of interactions with the average out-group
$p_{2}=f_{1|0g}+rp_{gi}.$ (S.21)
Then, the log-fitness can be written compactly as
$\frac{1}{K}\ln w_{gi}=a+bf_{1|0g}+[rb+df_{1|1g}-c]p_{gi}.$ (S.22)
This equation is again similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$. The interesting new possibility is that the agents’ behaviors are now different within and across groups in the sense that reciprocity $f_{1g^{\prime}|1g}<f_{1g|1g}$, and exploitation $f_{1g^{\prime}|0g}>f_{1g|0g}$, for $g^{\prime}\neq g$. This can lead to a variety of outcomes. However, to finish the analysis here, we consider the situation when states across groups are statistically independent. Then $r_{g^{\prime}}=0$ for $g^{\prime}\neq g$. This leads to two different contributions to the log-fitness
$\frac{1}{K}\ln w_{gi}=a+\underbrace{\frac{k_{g}}{K}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}+\underbrace{\sum_{g^{\prime}\neq g}\frac{k_{g^{\prime}}}{K}\left[\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\right]}_{\rm out-groups}.$ (S.23)
This presents a conundrum: while in-group contribution drives cooperation, out-group interactions select for defection. In general, the out-group is larger, so cooperation may be disrupted by population interference in group dynamics. There are two ways out of this dilemma. To better appreciate this, we use the simplifying notation
$\sum_{g^{\prime}\neq g}\frac{k_{g^{\prime}}}{K}p_{g^{\prime}}=(1-\frac{k_{g}}{K})p_{\bar{g}}.$ (S.24)
where $p_{\bar{g}}$ is the probability that an agent in the out-group cooperates. Note that this leads to a simpler expression
$\frac{1}{K}\ln w_{gi}=a+\frac{k_{g}}{K}\left(r_{g}b+f_{1g|1g}d-p_{\bar{g}}d\right)p_{gi}+p_{\bar{g}}dp_{gi}-p_{gi}c+(1-\frac{k_{g}}{K})p_{\bar{g}}b,$ (S.25)
so that the condition for cooperation to emerge is now
$\frac{k_{g}}{K}\left(r_{g}b+f_{1g|1g}d-p_{\bar{g}}d\right)+p_{\bar{g}}d>c,$ (S.26)
as shown in Table 1D.
We see that there are two different paths to cooperation. The first is that the agent interacts only within its in-group $k_{g}/K\rightarrow 1$. But this foregoes fitness benefits from interactions with out-group members. The second way out is to cooperate or defect conditionally on the group identity of the interaction partner. Then the in-group and out-group $p_{gi|g^{\prime}}$ can adapt differentially so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$. This may still mean that the average cooperative behavior $p_{gi}=\frac{k_{g}}{K}p(gi|g)+\frac{K-k_{g}}{K}p(gi|g^{\prime}\neq g)$ is small, if most interactions are external, but cooperation will now have acquired structure that is associated with information about group identity. Gradually extending cooperative behavior to selective out-groups also provides a pathway for the growth of cooperation aided by recognizable collective signals, leading to fluid structures of intergroup cooperation and competition as shown in Figure 3.
Supplementary Figures
Figure S1: A guide to interpreting Fig.2B. The left diagram plots strategy boundaries parameterized by $f_{1|1},q$ along lines of constant $f_{1|0}$. Points above each trace define parameters that promote cooperation for that value of $f_{1|1}$. The shaded regions define points where cooperation or defection are promoted for all values of $f_{1|0}$. That is, the asymptotic strategies are independent of the likelihood of exploitation. In the white region, the asymptotic strategy depends on which value of $f_{1|0}$ at which the game is evaluated. The right figure illustrates these conditions along lines of constant $q$. To demonstrate how these figures complement one another, we have highlighted the results of the plot for $q=5/12$, and the slice this corresponds to on the left plot. The boundaries of the strategy zones independent of $f_{1|0}$ are matched up with dashed lines that lead to arrows, and lie entirely outside of the range of the $q=5/12$ trace. We show that in the white region, $f_{1|1},f_{1|0}$ share a linear relationship.

### v2

@@
 We also show how this approach leads to powerful tools to analyze challenging problems of evolution in variable environments, game theory, and selection in group-structured populations.
-The approach is general and creates an explicit bridge between population dynamics, statistical learning theory, and emerging models of artificial intelligence.
+The approach is general and creates an explicit formal bridge between population dynamics, statistical learning theory, and emerging models of artificial intelligence.
 keywords:
@@
 Introduction
-Evolutionary modeling and theory have advanced rapidly over the past century incorporating the theory of replicators [1, 2, 3, 4], sources of intragenomic variation [5, 6, 7], environmental mechanisms of genetic variation [8, 9], and socio-ecological interactions via game theory [10] and the Price equation[11].
+Evolutionary modeling and theory have advanced rapidly over the past century incorporating the theory of replicators [1, 2, 3, 4], sources of intragenomic variation [5, 6], environmental mechanisms of genetic variation [7, 8], and socio-ecological interactions via game theory [9] and the Price equation[10].
 All of these approaches rely on the concept of fitness.
@@
 Higher relative fitness of a type results in higher prevalence in the population over time.
-Most studies of evolutionary dynamics adopt this framework to identify better-adapted phenotypes in natural settings in theoretical models [13, 14, 15] and laboratory studies of gene expression  [16].
-However, it was pointed out that this definition of fitness is incomplete and logically circular.
+Most studies of evolutionary dynamics adopt this framework to identify better-adapted phenotypes in ecosystems, theoretical models [12, 13, 14], and laboratory studies of gene expression  [15].
+However, it has been pointed out that this definition of fitness is incomplete and logically circular.
 This is because it does not explicitly refer to environmental adaptation [16, 17].
-A related issue deals with the nature of adaptive capacity, which must be a persistent property of phenotypes in variable environments and cannot depend on single events [16, 19].
-Thus, any solution to this problem must include how fitness is averaged over time and environmental states.
+This omission leads to a number of interrelated problems.
+The first problem deals with the nature of adaptive capacity, which must be a persistent property of phenotypes in variable environments and cannot depend on single events [15, 18].
+Any solution to this problem must explain how fitness emerges from the accumulation of exposures over time and environmental states.
 This issue also surfaces when modeling strategic interactions, such as in evolutionary game theory [19, 20].
-Finally, and related to temporal and population averaging, is the question of how fitness combines across scales.
-It is well known in replicator models that contributions to fitness multiply [12].
-This raises questions about the approximations to fitness as a linear function of traits in both modeling and empirical analysis [22, 23].
+Related to temporal and population averaging, is the crucial question of how fitness combines across scales.
+It is well known in replicator models that contributions to fitness multiply [1, 11, 21].
+This raises questions about approximations to fitness as a linear function of traits in both modeling and empirical analysis [22, 23].
 The linearity of contributions to fitness is often justified in situations of weak selection.
 This is commonly assumed in the analysis of kin and group selection [22, 24] and in gene expression experiments [25].
-Because many results and methods are based on the linearity of fitness, it becomes important to place this assumption on a stronger basis and to consider circumstances when selection is not weak [26, 27, 28].
+Because many results and methods are based on the linearity of fitness, it becomes important to place this assumption on a stronger basis and develop methods that are suitable for circumstances when selection is not weak  [26, 27, 28, 29].
 Other problems arise from explicit environmental or social selection, for example, in the context of evolutionary game theory.
-An illustration is the difficulty of deriving definite predictions from strategic interaction games for the problem of cooperation or altruism [29, 24, 30].
+An illustration is the difficulty of deriving definite predictions from strategic interaction games for the evolution of cooperation, or altruism [30, 24, 29].
 A related problem is the proliferation of Hamilton-type rules for similar models [31].
 These complications have led to the extensive use of computer simulations with different fitness definitions [32, 33], network and group structures [34, 35], and strategy spaces [36, 37, 38, 39].
-This body of work has emphasized repeated interactions [40, 41] and different types of reciprocity [31, 29] as fundamental conditions for cooperation to emerge and the need for group formation and intergroup competition [42, 22, 43].
-However, these findings remain contested [24], in part because of the difficulty of analyzing the impact of assumptions and modeling choices.
+This important body of work has emphasized repeated interactions [40, 41] and different types of reciprocity [31, 30] as fundamental conditions for cooperation to emerge and the need for group formation and intergroup competition [42, 22, 43].
+However, these findings remain contested and relatively disorganized [24], in part because of the difficulty of analyzing the impact of assumptions and modeling choices.
 Here, we show that these conceptual and practical problems share a common root in the theoretical foundations of fitness.
@@
 We take these ideas further by showing how they lead to the interpretation of fitness as a Bayesian likelihood, which is the optimal predictor of environmental states.
-We also show how fitness hence constructed allows us to build models of phenotype-environment selection in simple and clear ways in a variety of complex problems.
+We also show how fitness, hence constructed, allows us to build and analyze models of phenotype-environment selection in simple and clear ways for a variety of complex problems.
 This includes the effects of probabilistic environments in task selection, evolutionary dyadic games, and selection in group-structured populations.
-Phase spaces for these problems predict evolutionary transitions, expressed in the language of probability and information.
+Phase spaces of stable strategies for these problems predict evolutionary transitions, expressed in the language of probability and information.
 Perhaps most satisfying, this approach shows how selection in evolutionary dynamics is a form of optimal learning theory in populations and how fitness maximization corresponds to the fundamental principle of maximizing information between population structure and the statistics of ecologically relevant environments.
@@
 We now motivate how fitness quantifies adaptation to stochastic environments, and how this identification addresses the issues raised in the Introduction.
-Population replicators and their growth dynamics
+Population replicators and growth dynamics
 We define a structured population evolving under replicator dynamics.
-The population is made up of different types, $s$, such that the total population size $N=\sum_{s}N_{s}$, where $N_{s}$ is the population of each type at a given time.
-The temporal evolution follows the standard replicator equation for a haploid population in discrete time $\tau$,
-$N_{s}(t)=w_{s}N_{s}(t-\tau),\quad{\rm so\ that}\qquad N(t)=\sum_{s}N_{s}(t)=\sum_{s}w_{s}N_{s}(t-\tau)=wN(t-\tau),$ (1)
+The population is made up of different types, $s$, such that the total population size, $N=\sum_{s}N_{s}$, where $N_{s}$ is the population of each type at a given time.
+The temporal evolution follows the standard replicator equation for a haploid population in discrete time $\tau$, $N_{s}(t)=w_{s}N_{s}(t-\tau),$ so that
+$N(t)=\sum_{s}N_{s}(t)=\sum_{s}w_{s}N_{s}(t-\tau)=wN(t-\tau),$ (1)
 where $w_{s}\geq 0$ is the type’s fitness.
@@
 The growth rate, $\gamma_{s}$, is related to fitness by exponentiation $w_{s}=e^{\gamma_{s}\tau}\simeq 1+\gamma_{s}\tau$, and, correspondingly $\gamma_{s}=\frac{1}{\tau}\ln w_{s}\simeq\frac{w_{s}-1}{\tau}$.
-These relationships apply in continuous time while the approximations are first-order series expansions in $\tau$, Eq.
-1.
+These relationships apply in continuous time while the approximations are first-order series expansions.
 In the following, we set $\tau=1$ for simplicity.
@@
 However, Eq.
-1 and the standard set up of population genetics [1] make no explicit reference to $e$ We now show how these perspectives can be reconciled .
+1 and the standard set-up of population genetics [1] make no explicit reference to $e$: We now show how these perspectives can be reconciled.
 In evolutionary dynamics, the focus of optimization is the maximization of relative fitness. The time evolution of type frequencies follows from Eq. 1 as
-$f_{s}(t)=\frac{N_{s}(t)}{N(t)}=\frac{w_{s}}{w}\frac{N_{s}(t-1)}{N(t-1}=\frac{w_{s}}{w}f_{s}(t-1)\qquad\rightarrow\qquad p(H|D)=\frac{p(D|H)}{p(D)}p(H).$ (2)
-A number of authors have recently noted that the replicator equation under pure selection maps exactly to Bayes relation [47, 46], written on the right side of Eq.
-2.
+$f_{s}(t)=\frac{N_{s}(t)}{N(t)}=\frac{w_{s}}{w}\frac{N_{s}(t-1)}{N(t-1)}=\frac{w_{s}}{w}f_{s}(t-1).$ (2)
+The mapping to Bayes’ theorem follows as
+$f_{s}(t)=\frac{w_{s}}{w}f_{s}(t-1)\rightarrow p(H|D)=\frac{p(D|H)}{p(D)}p(H).$ (3)
+Several authors have recently noted that the replicator equation under pure selection, Eq.
+2, maps exactly to Bayes relation [47, 46], written on the right side of Eq.
+3.
 Bayes theorem gives the exact update of an initial probability that a hypothesis $h\in H$ is true, $p(h)$ (the prior), given new evidence $D$.
-This is written as the updated (posterior) probability $p(H|D)$.
-Because Bayes’ theorem is an identity that follows from the rules of probability, its procedure is the optimal way to incorporate new evidence to update the initial probability [48].
-As $p(H|D)$ is more constrained than $p(H)$ the data effectively selects specific hypotheses as more likely in a process where information was learned about the hypothesis space.
+This is written as the updated (posterior) probability $p(H|D)$ over all $h\in H$.
+Because Bayes’ theorem is an identity that follows from the rules of probability, this rule is the optimal way to incorporate new evidence to update any initial probability [48].
+As $p(H|D)$ is more constrained than $p(H)$ the data effectively selects specific hypotheses as more likely in a process where information is learned from the data about the hypothesis space.
 The key quantity in this procedure is $p(D|H)$, known as the likelihood.
@@
 To see this, let us map the data, $D$, to environmental states, $D\rightarrow e$, and hypotheses to population types, $H\rightarrow s$.
-We then identify $f_{s}(t)=f(s|e)$, where $e$ is the state of the environment at time $t$.
-$f(s|e)=\frac{p(e|s)}{{\hat{p}}(e)}f(s),\rightarrow w_{s}(e)=A(e)p(e|s),\quad w(e)=\sum_{s}w_{s}(e)f(s)=A(e)\sum_{s}p(e|s)f(s)=A(e){\hat{p}}(e).$ (3)
-This leads to the identification of fitness with the Bayesian likelihood, $w_{s}(e)=A(e)p(e|s)$, up to a multiplicative constant $A(e)$.
+We then identify $f_{s}(t)=f(s|e)$, where $e$ is the state of the environment at time $t$, leading to
+$f(s|e)=\frac{p(e|s)}{{\hat{p}}(e)}f(s)\rightarrow w_{s}(e)=A(e)p(e|s),\quad$ (4)
+where the fitness can be expressed in terms of the likelihood,
+$w(e)=\sum_{s}w_{s}(e)f(s)=A(e)\sum_{s}p(e|s)f(s)=A(e){\hat{p}}(e).$ (5)
+This leads to the identification of fitness with the Bayesian likelihood, $w_{s}(e)=A(e)p(e|s)$, up to a multiplicative constant in the phenotypes, $A(e)$.
 The factor of $A(e)$, is necessary because fitness, defined in Eq.
@@
 The factor $A$ disappears in relative fitness, but contributes to absolute fitness and to the growth rate of the entire population.
-Note that ${\hat{p}}(e)$ that appears in the average fitness is the estimate of the true environmental probability, $p(e)$, obtained by averaging the likelihood across types.
-This achieves a definition of fitness for each type $s$ as a variable probabilistic model of the environment, addressing the circularity problem.
-The picture of adaptation that emerges is explicitly relational and statistical: Selection relies on each phenotype having a generative model of the environment $p(e|s)$ and it is the goodness of this model, suitably averaged, that will translate into reproductive success.
+Note that ${\hat{p}}(e)$ in the average fitness is the estimate of the true environmental probability, $p(e)$, obtained by averaging the likelihoods across types.
+This achieves a definition of fitness for each type $s$ as a variable probabilistic model of the environment addressing the circularity problem.
+The picture of adaptation that emerges is explicitly relational and statistical: Selection relies on diverse phenotypes as alternative generative models of the environment, $p(e|s)$.
+The goodness of these models versus the statistics of events in the environment over time results in relative reproductive success.
 Fitness, growth rates, and information
 The fitness defined in Eq.
-3 is a stochastic quantity because it is evaluated at (random) states of the environment.
-Consequently, we must discuss how to average it over the environment and what quantities are maximized by the population dynamics of Eq.
+5 is a stochastic quantity because it is evaluated at successive (random) states of the environment.
+Consequently, we must discuss how to average it and what quantities are maximized by the population dynamics of Eq.
 2.
@@
 $\ln w_{s}(e)=\ln A(e)+\ln p(e)+\ln\frac{p(e|s)}{p(e)},$ (6)
-which is the instantaneous growth rate for type $s$.
-The first term shows the payoffs from the environment, the second measures environmental variability.
-The last term measures how close the $s$-type model of the environment is to the real thing, $p(e)$.
-Only the last term depends on the type and contributes to relative fitness.
-When the population dynamics is iterated, the types with the highest average log-fitness are positively amplified. This emerges from the temporal dynamics because
-$w_{s}(E)=\prod_{t=0}^{T}A(e_{t})p(e_{t}|s),\quad\rightarrow\ {\bar{\gamma}}_{s}=\sum_{i=1}^{T}\ln A(e_{t})p(e_{t}|s)\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (5)
-Here, ${\bar{\gamma}}_{s}$ is the average growth rate of type $s$ over a long time $T$ and $\ln A$ is the average of $\ln A(e)$, see Supplementary Information.
-The long-term temporal average becomes the average over $e$ due to the law of large numbers.
+which is the instantaneous growth rate, $\gamma_{s}$.
+The first term shows the payoffs from the environment, the second measures environmental variability, the third term measures how close the $s$-type model of the environment is to the real thing, $p(e)$, as we show next.
+Only this last term depends on the type and contributes to relative fitness.
+When the population dynamics is iterated over a long time $T$, the types with the highest average log-fitness over such period are positively amplified.
+This emerges from the temporal dynamics over $T$ because
+$w_{s}(E)=\prod_{t=0}^{T}A(e_{t})p(e_{t}|s)\rightarrow\ {\bar{\gamma}}_{s}=\frac{1}{T}\sum_{i=1}^{T}\ln A(e_{t})p(e_{t}|s).$ (7)
+It follows that
+${\bar{\gamma}}_{s}\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (8)
+Here, ${\bar{\gamma}}_{s}$ is the average growth rate of type $s$ over a long temporal history.
+The approximation converges to the actual rate because the long-term temporal trajectory effectively averages over $e$ due to the law of large numbers.
+All three terms have interesting meanings.
+The quantity $\ln A$ is the average of $\ln A(e)$, see Supplementary Information.
+It is a property of the environment, setting the magnitude of the largest possible growth rate.
+The second term is recognizable as the negative Shannon entropy of the environment, $-H(E)=\sum_{e}p(e)\ln p(e)$: it is always negative and larger in more variable environments, which therefore entail a growth penalty because they are more difficult to predict.
+The last term is $-D(p(E)||p(E|s)\leq 0$, which is the negative relative entropy, or Kullback-Leibler divergence [50].
+This quantity measures the distance in probability between the two distributions, meaning that it is maximized for the type $s$ with the perfect (or best) model of the environment $p(e|s)\rightarrow p(e)$.
+Other types will therefore incur a relatively smaller growth rate.
 Thus, the population dynamics selects the types $s$ that maximize the average growth rate, which can be interpreted in terms of information quantities [51, 52].
-The relative growth rate is maximized for $s$ such that the last term (which is the relative entropy and is negative definite) vanishes.
-This happens when the type’s fitness $p(e|s)\rightarrow p(e)$ becomes a perfect model of the real environment.
-The average of the type’s growth rates over the posterior $f(s|e)$, results in the mutual information $I(E,S)$, see Supplementary Information.
-In this sense, selection (1-3) converges on the type distribution that maximizes the information between population structure and the statistics of the environment.
+The average of the type’s growth rates over the posterior $f(s|e)$, results is the mutual information $I(E,S)$ between the population distribution over types and the probability of environmental states, see Supplementary Information.
+In this sense, selection (1-5) converges on the type distribution that maximizes the information between population structure and the statistics of the environment.
 Two additional consequences follow from the definition of fitness as a Bayesian likelihood.
 First, because fitness is a probability, statistically independent contributions combine as products.
-This means that fitness will only rarely be a linear combination of factors as is often assumed.
-Second, independent contributions do add up in the log-fitness, as they do in all information quantities.
-This makes the average log-fitness (or growth rate) the natural quantity for the analysis of complex evolutionary problems, even away from weak selection.
+This means that fitness will only rarely be a linear combination of factors, as often assumed.
+Second, independent contributions do add up in the log-fitness, as in all information quantities.
+This makes the average log-fitness (growth rate) the natural quantity for analyzing complex evolutionary problems, especially away from weak selection.
 Fitness in probabilistic environments
@@
 For simplicity, we consider two tasks, denoted $s=0,1$, in a binary environment, $e=0,1$.
-The standard setup prescribes rewards, $Y_{s}(e)$, to each of these state combinations, Table 1A-B.
+The standard setup prescribes rewards, $Y_{s}(e)$, to each state combination, Table 1A-B.
 To understand the dynamics of selection, note that it pays off to align the task with the environment $s=e$, regardless of the specific state, $e=0,1$ because $Y_{s=e}=Y_{+}>Y_{s\neq e}=Y_{-}$.
@@
 The identification of population dynamics of this type with iterated Bayesian selection requires mapping the rewards $Y_{s}(e)$ to the likelihood function, $p(e|s)$.
-Because the likelihood is a conditional probability model, it is positive and normalized, meaning that any mapping to payoffs must satisfy such requirements.
+Because the likelihood is a conditional probability, it is positive and normalized, meaning that any mapping to payoffs must satisfy such requirements.
 We want mappings that preserve the relative strength of the payoffs.
@@
 We can write the fitness as
-$w_{s}(e)=Y_{+}^{(1-s)(1-e)+se}Y_{-}^{(1-s)e+s(1-e)},\qquad\ln w_{s}=\ln Y_{+}+\ln\frac{Y_{+}}{Y_{-}}\left(2se-s-e\right).$ (6)
-Note that we can now obtain the probability of $s$ that will maximize fitness over a long (stationary) history of the environment by taking expected values over $e$ and $s$.
+$w_{s}(e)=Y_{+}^{(1-s)(1-e)+se}Y_{-}^{(1-s)e+s(1-e)}.$ (9)
+The logarithm of the fitness (log-fitness) is then
+$\ln w_{s}=\ln Y_{+}+\ln\frac{Y_{+}}{Y_{-}}\left(2se-s-e\right).$ (10)
+We can now obtain the probability of $s$ that maximizes fitness over a long (stationary) history of the environment by taking expected values over $e$ and $s$.
 We define $f(s=1)=p$ and $p(e=1)=q$, and $f_{1|1}=f(s=1|e=1)$ and $f_{1|0}=f(s=1|e=0)$.
 The log-fitness is maximized when the quantity $(2f_{1|1}-1)q-p$ is maximal.
-This requires $f_{1|1}=1$, which implies $p=f_{1|1}q+f_{1|0}(1-q)=q$, which means that the population distribution of tasks is equal to the environmental probability.
-We can also write the dynamics of selection by iterating the likelihood, given an environmental history and an initial distribution of traits $s$ in the population, see Supplementary information
+This requires $f_{1|1}=1$, which implies $p=f_{1|1}q+f_{1|0}(1-q)=q$.
+Hence, the population task distribution evolves to mirror the environmental probability.
+We can also write the dynamics of selection by iterating the likelihood, given an environmental history and an initial distribution of traits $s$, see Supplementary Information
 $f(s|e)\sim\Pi_{t=1}^{T}p(e_{t}|s_{t})f(s)=Y_{+}^{N_{e=s}}Y_{-}^{T-N_{e=s}}f(s),$ (11)
-where $N_{e=s}$ is the number of times that $s=e$.
-For large times, the law of large numbers tells us that $N_{e=s}\rightarrow Tp(e,s)=Tp(s|e)p(e)$.
+where $N_{e=s}$ is the number of instances that $s=e$ over $T$ times.
+For long times, the law of large numbers tells us that $N_{e=s}\rightarrow Tp(e,s)=Tp(s|e)p(e)$.
 The likelihood is maximized by taking $p(s=e|e)\rightarrow 1$, which means that $f(s)\rightarrow p(e)$, as derived above.
@@
 The population now becomes an environment from which each agent derives its fitness.
-This contrasts with the task selection problem, where environmental dynamics are stationary and therefore independent of agent actions.
-The simplest situation is a mixture of two games: One game when the state of the environment $e=0$, and another when $e=1$.
-The total payoff over time will then be the result of agents recognizing which environmental situation they are in and adjusting their interactions accordingly.
-This also generalizes the task-switching problem to what we may call a game-switching situation.
-We first demonstrate how the statistical treatment of fitness simplifies the analyses of dyadic games.
-Consider first the prisoners’ dilemma (PD), used for decades to analyze the evolution of cooperation [29, 20].
-Table 1C-D shows the game’s payoff structure, where $s=0$ corresponds to defection and $s=1$ to cooperation.
-We consider two players $s_{1},s_{2}$ each with probability cooperation $p_{1},p_{2}$, respectively.
-Payoffs are chosen such that $W>R>L>P>0$.
+This contrasts with the task selection problem, where environmental dynamics are stationary and therefore independent of the agent’s actions.
+We first demonstrate how the statistical treatment of fitness simplifies the analysis of dyadic games.
+Consider the prisoners’ dilemma (PD), used for decades to analyze the evolution of cooperation [30, 19].
+Table 1C-D shows the game’s payoff structure, where $s=0$ denotes defection and $s=1$ cooperation.
+We consider two players $s_{1},s_{2}$ each with cooperation probability $p_{1},p_{2}$, respectively.
+Payoffs are such that $W>R>L>P>0$.
 We write $s_{1}$’s fitness as
@@
 Eq.
-8 gives stochastic pay-offs at each interaction rather than on average, which would requires additional assumptions [52, 53].
+12 gives stochastic pay-offs at each interaction rather than on average, which would require additional assumptions [53, 54].
 This leads to
-$\ln w_{s_{1}}=a-cs_{1}+bs_{2}+ds_{1}s_{2},\qquad a=\ln AL,\ c=\ln\frac{L}{P},\ b=\ln\frac{W}{L},\ d=\ln\frac{RL}{PW}.$ (9)
-This form of log-fitness coincides with the typical parameterization of fitness as a linear function of costs and benefits of cooperation in the classical problem of altruism.
-Specifically, there is a fitness cost $c>0$ of cooperating when $s_{1}=1$, an independent benefit $b>0$ from the cooperation of others, $s_{2}=1$.
+$\ln w_{s_{1}}=a-cs_{1}+bs_{2}+ds_{1}s_{2},$ (13)
+where $a=\ln AL$, $c=\ln\frac{L}{P}$, $b=\ln\frac{W}{L}$, and $d=\ln\frac{RL}{PW}$.
+This polynomial log-fitness coincides with the typical parameterization of fitness as a linear function of costs and benefits of cooperation in the classical problem of altruism [30, 19].
+Specifically, there is a fitness cost, $c>0$, of cooperating, and an independent benefit $b>0$ from the cooperation of others.
 There is also an extra benefit (if $d>0$) when both agents cooperate, which was considered by Queller [23, 22].
 Note also that $d=c-\delta<c$, with $\delta=\ln W/R>0$.
-We see that, in a single-shot game, $s_{1}=0$ (defection) results in higher log-fitness than cooperation, even when the other player cooperates ($s_{2}=1$).
-This is the familiar Nash equilibrium, because the log-fitness of $s_{2}$ is identical with labels interchanged.
+In a single-shot game, $s_{1}=0$ (defection) results in higher log-fitness even when the other player cooperates ($s_{2}=1$).
+This is the game’s familiar Nash equilibrium, because the log-fitness of $s_{2}$ is identical with labels interchanged.
 A Fitness, $w_{s}(e)$ $e=0$ $e=1$ $s=0$ $Ap(0|0)$ $Ap(1|0)$ $s=1$ $Ap(0|1)$ $Ap(1|1)$ B Returns, $Y_{s}(e)$ $e=0$ $e=1$ $s=0$ $Y_{+}$ $Y_{-}$ $s=1$ $Y_{-}$ $Y_{+}$ C $Y_{s_{1},s_{2},e=0}$ $s_{2}=0$ $s_{2}=1$ $s_{1}=0$ $L,L$ $W,P$ $s_{1}=1$ $P,W$ $R,R$ D $Y_{s_{1},s_{2},e=1}$ $s_{2}=0$ $s_{2}=1$ $s_{1}=0$ $L,L$ $P,W$ $s_{1}=1$ $W,P$ $R,R$
@@
 E: Summary of conditions for the emergence of cooperation.
-By contrast, the goal of evolutionary game theory is the identification of selected strategies over long histories of interaction, which is analyzed via averaging Eq.
-9.
+By contrast, the goal of evolutionary game theory is the identification of selected strategies over long histories, which is analyzed via averaging Eq.
+13.
 There are three distinct situations with growing levels of complexity.
@@
 And third, when their phenotypes become (statistically) dependent.
-Each of these general situations is associated with different population outcomes.
+Each of these general situations is associated with different outcomes.
 When the agents are statistically independent with cooperation probabilities $p_{1},p_{2}$, the average log-fitness is
@@
 In this sense, it must correspond to the single shot game and has the same optimal solution.
-Note also that, because of statistical independence, there can be no information between the agents.
+Note also that statistical independence means that there cannot be any information between the agents.
 Now consider a seemingly small but important change in these assumptions.
 We consider $p_{1}=p_{2}=p$, which is frequently assumed in population genetics.
-It is dubbed random mating (in diploid populations) or a fully mixing population [1].
-This results in
-$\ln w_{s_{1}}=a-(c-b)p+dp^{2},$ (11)
-The polynomial form of the log-fitness in $p$ is familiar and simple to analyze.
+This is often dubbed random mating (in diploid populations) or a fully mixing population [1], resulting in
+$\ln w_{s_{1}}=a-(c-b)p+dp^{2}.$ (15)
+The quadratic form of the log-fitness in $p$ is familiar and simple to analyze.
 A distinction emerges depending on the relative magnitude of $b$ and $c$, Fig.
@@
 When $b>c$ (or $WP>L^{2}$) the situation is simple because the log-fitness is always a growing function of $p$, with a single maximum at $p=1$ corresponding to full cooperation.
-Whatever the initial state of the population, cooperation will always emerge as the dominant strategy and will eventually fixate.
+Thus, whatever the initial state of the population, cooperation will always emerge as the dominant strategy and will eventually fixate.
 Relative to the statistically independent case, cooperation now emerges from the fact that the population is effectively playing itself, so that increases in mutual exploitation lead to overall lower fitness.
@@
 The existence of this fitness minimum creates a local fitness maximum at $p=0$ (mutual defection), which is locally stable.
-In these circumstances, sufficient cooperation $p>p_{*}$ must be present for the population to converge to the global fitness maximum of full cooperation ($p=1$).
+In these circumstances, sufficient cooperation $p>p_{*}$ must be present for the population to converge to the global fitness maximum of full cooperation ($p=1$), Figure 2A.
 This corresponds to a simple condition for cooperation to emerge, related to Hamilton’s rule [55].
 From Eq.
-11, we identify benefits of cooperation as $B=b+dp$ and costs with $C=c$.
+15, we identify benefits of cooperation as $B=b+dp$ and costs with $C=c$.
 Then $B>C$ is a condition for full cooperation to emerge, which is equivalent to $p>p_{*}$.
-The population dynamics around $p=p_{*}$ is characterized by a first-order transition – a "tipping point"– because there is a valley of fitness to be crossed to reach the global maximum, starting with generalized defection.
-This can be done locally, over space or on networks, as many authors have noted based on numerical experiments [27, 55].
+The population dynamics around $p=p_{*}$ is characterized by a first-order transition, a "tipping point," because there is a valley of fitness to be crossed to reach the global maximum.
+This can be done locally, over space or on networks, as many authors have noted based on numerical experiments [27, 56, 38].
 We will return to this point below.
-Third, the situation is both more interesting and more complex when the agents’ actions become statistically dependent.
+Third, the situation is more interesting and more complex when the agents’ actions become statistically dependent.
 In this case, there are three parameters at play in the joint distribution: four joint probabilities minus the normalization.
 We can parameterize this space in analogy to task selection as $p_{2}=f_{1|0}(1-p_{1})+f_{1|1}p_{1}$ and $f(s_{1}=1,s_{2}=1)=f_{1|1}p_{1}$, with the shorthand $f_{1|1}=f(s_{2}=1|s_{1}=1)$, $f_{1|0}=f(s_{2}=1|s_{1}=0)$.
-Note that $r=f_{1|1}-f_{1|0}$ is the regression of $p_{2}$ on $p_{1}$, which vanishes when the two agents are statistically independent.
-Thus, $r$ is a measure of how (co-)related the two states are.
+Note that $r=f_{1|1}-f_{1|0}$ is the regression [10, 22] of $p_{2}$ on $p_{1}$, which vanishes when the two agents are statistically independent.
+Thus, $r$ is a measure of how (co-)related the two agents’ phenotypes are.
 The log-fitness becomes
@@
 This transition has a geometric interpretation in the reciprocity-exploitation ($f_{1|1},f_{1|0}$) plane.
-We can write the cooperation condition as a line, $y=f_{1|1}$, $x=f_{1|0}$, $y=y_{0}+\beta x$, with $y_{0}=\frac{c}{b+d}$, $\beta=\frac{b}{b+d}$ , as shown in Figure 2C.
+We can write the cooperation condition as a straight line, $y=f_{1|1}$, $x=f_{1|0}$, $y=y_{0}+\beta x$, with $y_{0}=\frac{c}{b+d}$, $\beta=\frac{b}{b+d}$ , as shown in Figure 2C.
 Cooperation, $p_{1}\rightarrow 1$, emerges for values above this line.
-To be above the line requires controlling exploitation, optimized when $f_{1|0}=f(s_{2}=1|s_{1}=0)\rightarrow 0$ (which implies that punishment or retaliation, $f(s_{2}=0|s_{1}=0)\rightarrow 1$.
-It also requires reciprocity meaning that $f_{1|1}=f(s_{2}=1|s_{1}=1)\rightarrow 1$ ( so that one becomes safe from exploitation, $f(s_{2}=0|s_{1}=1)\rightarrow 0$).
+To be above the line requires controlling exploitation, optimized when $f_{1|0}=f(s_{2}=1|s_{1}=0)\rightarrow 0$, which implies that punishment or retaliation is strong, $f(s_{2}=0|s_{1}=0)\rightarrow 1$.
+It also requires reciprocity meaning that $f_{1|1}=f(s_{2}=1|s_{1}=1)\rightarrow 1$, so that one becomes safe from exploitation, $f(s_{2}=0|s_{1}=1)\rightarrow 0$.
 Note that the first of these conditions may take the population towards mutual defection [57], which is suboptimal.
@@
 Below the line, cooperation cannot emerge.
-These outcomes agree with established results on the effects of punishment and rewards on cooperation [57, 58].
-That is, both punishment and reciprocity promote the establishment of cooperation but alone punishment does not suffice [59].
+These outcomes agree with established results on the effects of reciprocity, punishment and rewards on cooperation [58, 59].
+That is, both punishment and reciprocity promote the establishment of cooperation but alone punishment does not suffice [60, 57].
 The information in the population is maximal as agents coordinate but the amount of information at stake converges to zero at full cooperation because then there is no uncertainty left to explain.
@@
 Cooperation emerges in each case for parameters above the lines.
-See also Fig.S1
+See also Fig.
+1 of the supplement
 Mixed games in variable environments
@@
 Selected strategies depend on the relative probabilities of the two games set by the environment $q=p(e=0)$. The log-fitness becomes
-$\ln w_{s_{1}}=a-c(e)s_{1}+b(e)s_{2}+ds_{1}s_{2},\quad{\rm with}\quad c(e)=c(1-e)-be,\ b(e)=b(1-e)-ce.$ (13)
+$\ln w_{s_{1}}=a-c(e)s_{1}+b(e)s_{2}+ds_{1}s_{2},$ (17)
+with $c(e)=c(1-e)-be$ , $b(e)=b(1-e)-ce$.
 Costs and benefits swap between the PD and the ToC, for which $\ln w_{s_{1}}=a+bs_{1}-cs_{2}+ds_{1}s_{2}$.
 If the two agents act independently, then the agent should always cooperate in the ToC because both $b,d>0$.
-If the two agents act independently but share the same probability $p_{1}=p_{2}=p$, the log-fitness is independent of $e$ because there is a symmetry between the two agents so that the benefits and costs of cooperation are fully socialized.
+If the two agents share the same probability $p_{1}=p_{2}=p$, the log-fitness is independent of $e$ because there is a symmetry between the two agents so that the benefits and costs of cooperation are fully socialized.
 Thus, in the fully mixed population, changes in the environment have no effect on the average population dynamics and the phase diagram is the same as in Fig.2B.
 When the two agents coordinate, the situation becomes more interesting.
-If the agents can tell which game they are playing at each time, their probability of cooperation will be the selected strategy for each game weighed by their relative probability $q$.
-The other situation we consider is when agents act conditionally on each other independently of the state of the environment.
+If agents can tell which game they are playing at each time, their probability of cooperation will simply follow from the selected strategy for each game, weighed by their relative probability, $q$.
+Alternatively, we consider when agents act conditionally on each other, independently of the state of the environment.
 Then we obtain the same condition for cooperation to emerge as in PD, but with costs and benefits that depend on the environment as in Eq.
-13.
+17.
 Fig.
@@
 At $q=1/2$ the slope is zero.
-Eventually, at $q=2/3$, ToC is sufficiently dominant and cooperation becomes unconditional.
-This analysis is general and can be obtained for any combination of games leading to complex and interesting phase spaces, Fig.
+Eventually, at $q=2/3$, ToC is sufficiently dominant and cooperation emerges unconditionally.
+This analysis is general and can be obtained for any combination of games, leading to complex and interesting phase spaces, Fig.
 2.
-As a result, we obtain transitions between different (optimal) population states as functions of environment probabilities, which play the role of control parameters, analogous to temperature in physical systems.
+As a result, we obtain transitions between different (optimal) population states as functions of environment probabilities, which play the role of control parameters analogous to temperature or pressure in physical systems.
 We have therefore shown how mixing the prisoner’s dilemma with a cooperative game increases the selection for cooperation in non-trivial ways.
-This is likely a more realistic situation, as organisms and especially people, have a diversity of modes of interaction some competitive and some cooperative.
-Thus, the presence of explicitly cooperative channels some of the time can help fix unstable cooperation all of the time.
+This is likely a more realistic situation as organisms, and especially people, have a diversity of modes of interaction, some competitive and some cooperative.
+Thus, the presence of explicitly cooperative channels some of the time helps unstable cooperation to emerge all of the time.
 Selection in groups
-The probabilistic interpretation of fitness readily extends to structured populations.
+The probabilistic interpretation of fitness extends naturally to structured populations.
 Within a population, group identities specify additional layers of organization that condition phenotypes and therefore may contribute to individual selection.
 We write the fitness of individual $i$ in group $g$ as $w_{gi}(g^{\prime},j)=Ap(\{s_{g^{\prime}j}\}|s_{gi})$, where $\{s_{g^{\prime}j}\}$ denotes the joint state of all neighbors $j,g^{\prime}$.
-Typically, these neighbors are not statistically independent from each other leading to synergistic interactions[23, 60].
+Typically, these neighbors are not statistically independent, leading to synergistic interactions [23, 61].
 This is important for finding clusters of states (genes, phenotypes) that collectively encapsulate adaptive functions [62].
-However, the most common situation arises when interactions are dyadic (between pairs, only), so that the game occurs on a simple graph [62, 63, 35, 64].
-In the log-fitness, this approximation is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that
-$w_{gi}=\Pi_{j,g^{\prime}}Ap(s_{g^{\prime}j}|s_{gi}),\rightarrow\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (14)
+However, the most common situation arises when interactions are dyadic (between pairs only), so that the game occurs on a simple graph [63, 64, 35, 56].
+In the log-fitness, this approximation is obtained when all the $s_{jg^{\prime}}$ are conditionally independent given $s_{ig}$ such that $w_{gi}=\Pi_{j,g^{\prime}}Ap(s_{g^{\prime}j}|s_{gi})$, resulting in
+$\ln w_{g,i}=\ln A+\sum_{g^{\prime},j}\ln p(s_{g^{\prime}j}|s_{gi}).$ (18)
 With these assumptions, we can write the log-fitness to be identical to Eq.
-10, where the second player is replaced with the average over all connections, $s_{2}=1/K\sum_{g^{\prime},j}{s_{g^{\prime}j}}$.
-All our previous results still apply but require attention to group dependencies (see Supplementary Information for details).
+14, with the second player replaced by the average over all connections, $s_{2}=1/K\sum_{g^{\prime},j}{s_{g^{\prime}j}}$.
+All our previous results still apply but require attention to group dependencies, see Supplementary Information for details.
 Figure 3: Selection in structured populations depends on group sizes and treatment of out-groups.
@@
 15 still applies and a phase transition occurs at $p=p_{*}$, if $b<c$.
-When the states are not independent, the log-fitness is written compactly in terms of contributions from each interaction, weighed by the size of the population $1/K$,
+When the states are not independent, the log-fitness is written compactly in terms of contributions from each interaction, weighed by population size $1/K$,
 $\frac{1}{K}\ln w_{gi}=a+bf_{1|0g}+[rb+df_{1|1g}-c]p_{gi}.$ (19)
-This equation is similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$, where the additional $g^{\prime},g$ labels denote conditional dependence across groups.
-The absence of the label means average across groups.
+This equation is similar to what we obtained above, but with population definitions of $r$, $f_{1|1g},f_{1|0g}$, where the additional $g^{\prime},g$ labels denote conditional dependence across groups; the absence of the label means average across groups.
 The interesting possibility is that the agents’ behaviors are now different within and across groups in the sense that reciprocity $f_{1g^{\prime}|1g}<f_{1g|1g}$, and exploitation $f_{1g^{\prime}|0g}>f_{1g|0g}$, for $g^{\prime}\neq g$.
-Treating out group interactions as statistically independent leads to
-$\ln w_{gi}=a+\underbrace{k_{g}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}+\underbrace{\sum_{g^{\prime}\neq g}k_{g^{\prime}}\big{[}\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\big{]}}_{\rm out-groups}.$ (16)
+Treating out group interactions as independent leads to
+$\frac{1}{K}\ln w_{gi}=$ $a+\underbrace{\frac{k_{g}}{K}\left(br_{g}+df_{1g|1g}-c\right)}_{\rm in-group}p_{gi}$ (20)
+$+\underbrace{\sum_{g^{\prime}\neq g}\frac{k_{g^{\prime}}}{K}\big{[}\left(dp_{g^{\prime}}-c\right)p_{gi}+bp_{g^{\prime}}\big{]}}_{\rm out-groups}.$ (21)
 This presents an interesting conundrum: The in-group contribution drives cooperation, but the out-group selects for defection.
@@
 First, the agent may cease to interact with out-group members so that $k_{g^{\prime}\neq g}=0$, but this foregoes fitness benefits from the possible altruism of a large population.
-Fig.3A shows when these benefits promote cooperation, solving $rB>C$ for $k_{g}/K$ as a function of $p_{\bar{g}}$.
+Fig.3A shows when cooperation emerges, solving $rB>C$ for $k_{g}/K$ as a function of $p_{\bar{g}}$.
 As before, the contours define a boundary, above which more cooperation increases fitness.
-We also plot this solution in Fig3B for $f_{1g|1g}$ and $p_{\bar{g}}$, which can be compared to Fig.2B when $q=0$.
+We also plot this solution in Fig.
+3B for $f_{1g|1g}$ and $p_{\bar{g}}$, which is comparable to Fig.
+2B when $q=0$.
 Second, cooperation or defection could conditionally depend on the group identity of the interaction partner.
-Then the in-group and out-group probabilities of cooperation $p_{gi|g^{\prime}}$ adapt differentially so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$.
+Then the in-group and out-group probabilities of cooperation $p_{gi|g^{\prime}}$ evolve in opposite directions so that $p(gi|g)\rightarrow 1$, and $p(gi|g^{\prime}\neq g)\rightarrow 0$.
 This may still mean that the average cooperative behavior $p_{gi}=\frac{k_{g}}{K}p(gi|g)+\frac{K-k_{g}}{K}p(gi|g^{\prime}\neq g)$ is small if most interactions are external, as illustrated in Fig.3A, but cooperation will now be stabilized within each group.
@@
 Coordination and cooperation can continue to spread across groups selectively, as larger sets of group markers form new collective identities associated with in-group behavior through the mechanisms described in Table 1D.
-In this way, pedigree relations may be the most natural basis for in-group formation [22], but mutualistic strategic interactions can also spread if they become associated with recognizable characteristics and beneficial phenotypic behaviors.
+In this way, pedigree relations (kin selection) may often be the most natural basis for in-group formation [22], but mutualistic interactions can also spread if they become associated with recognizable collective signals and beneficial phenotypic behaviors.
 Discussion
-We have demonstrated how to resolve several conceptual and practical problems in evolutionary theory by modeling fitness as an explicit process of adaptation to statistical environments.
-This result arises from interpreting population type frequencies as probabilities and assuming that adaptation is an optimal encoding of environmental evidence through Bayesian inference–a standard mapping in evolutionary dynamics [47, 46].
+We demonstrated that defining fitness quantitatively as an explicit process of adaptation to statistical environments resolves several fundamental conceptual and practical problems in evolutionary theory.
+This re-definition arises necessarily from interpreting population type frequencies as probabilities and assuming that adaptation is an optimal encoding of environmental evidence through Bayesian learning–an argument advanced recently by several authors [47, 46].
 Writing fitness as a Bayesian likelihood clarifies the construction of mathematical models of adaptation in terms of conditional probability functions.
@@
 To this end, many authors have identified information theory as a natural language for biological phenomena ranging from the evolution of genome organization [65, 66], the emergence of structure in ecological networks [67], and fitness as it relates to environmental stochasticity [11, 13].
-By formalizing the environment as a learnable statistical model, our approach replaces the circular notion of fitness as reproductive success by a well-defined statistical quantity that is optimized through recursive inference and characterized by information.
-In introducing this mechanism, we open up several directions for future research.
+By formalizing the environment in terms of learnable statistics associated with payoffs to growth, our approach replaces the circular notion of fitness as reproductive success by a well-defined statistical quantity that is optimized through recursive inference and characterized by information.
+In introducing this connection, we hope to open up several directions for future research.
 First, defining fitness as a likelihood can guide both field observations and laboratory studies of evolutionary dynamics.
@@
 Practical applications in empirical settings also require identifying sources of variation not associated with selection.
-In standard evolutionary theory, mutations deviate population dynamics from pure Bayesian updating.
+As in standard evolutionary theory, mutations deviate population dynamics from pure Bayesian updating.
 Although fitness is still updated through the likelihood function, the dynamics now reflect a mutation-selection balance.
-Non-selective forces such as genetic drift and gene flow prevent the complete optimization of information transfer between the population and the environment, leading to incomplete adaptation and fixation of dominant types.
-These factors introduce parameters, such as the magnitude of fluctuations, to the analysis which will typically increase the evolvability of populations and enrich phase-space possibilities for evolutionary transitions.
-Many of these possibilities will lead to increased group heterogeneity.
+Other non-selective forces such as genetic drift and gene flow, prevent the complete optimization of information transfer between the population and the environment, leading to incomplete adaptation and fixation of dominant types.
+These factors introduce parameters, such as the strength of inter-group interactions, to the analysis which may increase the evolvability of populations and enrich phase-space possibilities for evolutionary transitions [68].
+Many of these possibilities will lead to increased intra-group heterogeneity.
 Our analysis of dyadic interactions shows that clusters of related phenotypes can promote cooperative behavior, introducing an information signal that can drive selection as argued in the general problem of major evolutionary transitions [69].
@@
 This allows collective interactions to be analyzed via the structure of conditional (in)dependence of states, as established in probabilistic graphical models [70].
-These results strengthen the intersection of evolutionary biology and statistical learning theory by integrating evolutionary dynamics with causal inference, information theory, and statistical learning, thereby paving the way for interdisciplinary applications in artificial intelligence.
+These results strengthen the intersection of evolutionary biology and statistical learning theory by integrating evolutionary dynamics with causal inference, information theory, and recent progress in machine learning.
 Methods
@@
 This is due to the law of large numbers
-${\bar{\gamma}}_{s}=\sum_{t=1}^{T}\ln\big{[}A(e_{t})p(e_{t}|s)\big{]}\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (S.2)
+${\bar{\gamma}}_{s}=\frac{1}{T}\sum_{t=1}^{T}\ln\big{[}A(e_{t})p(e_{t}|s)\big{]}\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (S.2)
 which is also the average log-fitness. We can write these averages in terms of information quantities, leading to